OpenStudy (anonymous):

(12x+6)-5=x+34 Do I distribute -1 or -5????

OpenStudy (amistre64):

neither

OpenStudy (amistre64):

that is a subtract 5; not a distribute -5

OpenStudy (anonymous):

I have to distribute something though.........

OpenStudy (amistre64):

why? you just feel distributivey lately?

OpenStudy (amistre64):

the (..) are nothing at the moment by eye clutter

OpenStudy (amistre64):

distribute a "1" thru them if need be

OpenStudy (anonymous):

I was not taught this at all.......

OpenStudy (amistre64):

not taught what?

OpenStudy (amistre64):

(12x+6) - 5 = -5 + (12x+6) = -5 +12x +6

OpenStudy (anonymous):

Just look at the problem like this 12x + 6 -5=x+34

OpenStudy (anonymous):

the point is unless that says \[(12x+6)(-5)\] the parentheses are unnecessary and it might as well just say \[12x+6-5\]

OpenStudy (anonymous):

Ok, thank you, now that it's verified by three people, I feel more comfortable with it now, thank you everyone!